INAF queue observing, 2015 Nov 07-Nov 08

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg on call
Instrument: LBC-MODS-LUCI

Summary

At the sunset the sky is clear, the humidity is 20%

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 8 hrs 50 m (~90%)

Weather time loss: 0 h (0%)

technical time loss 1 h 20 m (~10%)

Data Summary

Project_RUN Objects FilterSorted ascending ExpTime Notes
Skyflat V, I
Clementini PegasusIII, PiscesII B, V 0.3 ob from number 11 to 13, average seeing 1.3-1.5'', OUT of SPECS
Cusano AndXXI B, V 15min ob 1 and 2
Cusano AndXIX B, V 15min ob 1 and 2
Cusano And_field_1 B, V 0.5h  
Bellazzini H 104.4+04.1 g, r 0.2h completed
Antoniucci v350 cep J (ima) 0.1 h completed
Antoniucci xztau J (ima) 0.1 h completed
Phot Stand SA98 SVBg_rVi    
Castellano MACSJ0717.5 Uspec, V 36 min  
Antoniucci v350cep zJspec, HKspec 0.6h (spec) completed
Antoniucci tell v350cep   zJspec, HKspec
Antoniucci xztau zJspec, HKspec 0.6h (spec) completed
Antoniucci tell xztau   zJspec, HKspec
Antoniucci cal   zJspec, HKspec

Details

00:57 running SkyFlats for V-I filters

01:40 running DOFPIA, we then start trying to collimate

01:47 epar lbcrangebal
  • DIMM seeing 1.4''
  • starting with Pegasus11 (and Pisces11)
  • seeing 1.3''-1.4''

02:09 Pegasus12
  • seeing 1.15''
02:14 Pisces12
  • seeing 1.4''

02:20 Pegasus13
  • seeing 1.2-1.3''

02:26 Pisces13

02:32 Pegasus 14

02:39 we are out of specs for Clementini (average seeing 1.4''), so we slew to Cusano.

02:57 lost some time for problems with IRAF (lbcrangebal), now running AndXIX _1.ob
  • seeing 1.5''
03:07 AndXXI _1.ob

03:14 probable readout problem in one of the chips of the red channel; the image is blurred. We have to restart LBC.

03:26 starting again DOFPIA and copointing for Cusano

03:36 we are not able to copoint because lbcrangebal doesn't give the offset. We try killing and starting Iraf again. Now it works.

03:39 running AndXXI _1.ob again
  • seeing 1.5''

  • 03:51 running DOFPIA

03:55 running AndXIX _1.ob
  • 04:03 thee is no tracking in the blue channel, we are finishing this integration (that should be ok) and then we have to power-cycle LBC again
  • the blue image is not good, red image is ok
  • seeing 1.5-1.6''

04:10 DOFPIA and copointing again for Cusano

04:19 AndXXI _2.ob
  • seeing 1.3''

04:28 DOFPIA and AndXIX _2.ob
  • seeing 1.5''

04:45 And_field_1.ob
  • dofpia
  • 4:51 run science
  • 5:00 seeing variable 1''-1.3''

05:23 switch to MODS

05:47 the seeing is very bad and variable: 1.4''-2.0'' we can not do the Saracco program, the blue channel with MODS2 doesn't work, so we switch to luci2 to do the backup program of Antoniucci

06:25 acquisition target V350 Cep
  • v350_cep.xml GO luci2.20151108.0001-0007
  • 6:49 seeing 2.35''
  • 06:57 begin science luci2.0008-0015

7:25 imaging V350 Cep
  • v350_cep_im.xml luci2.0016-0021

7:35 telluric star for the target V350 Cep * v350_cep_tell.xml GO * acquisition luci2.0022-0024 * science luci2.0025-00.28 * seeing 2''-3''

7:55 slew to the program XZTau xz_tau.xml GO
  • acquisition luci2.0029-0031
  • 8:10 problem with the air compressor at the primary mirror. we stop.

8:35 the TO gives us the permission to acquire again but he have to collimate again.

08:44 xz_tau.xml GO
  • the seeing is 0.9'', we continue with the backup program because we want to be sure that the seeing is stable
  • acquisition luci2.0032-0036
  • 08:56 detector focus commit error
  • 09:08 start the science luci2.0037-0044
  • seeing 0.8'''

09:14 run imaging: xz_tau_im.xml GO luci2.0045-0050

09:20 run the telluric star for this target: xz_tau_tell.xml GO
  • 09:24 lost the guide star
  • 09:26 start again
  • acquisition luci2.0051-0050-54 erro in the acquisition, luci2 takes two images of the slit but any of the source
  • the AS reinitialize and restart the scheduler we re-load start, but error during commit.
  • 09:38 The SA start the scheduler again, we re-load the script and go
  • acquisition luci2.0055
  • science luci2.0056-0059

09:52 the seeing is good, about 1.0'', switch to LBC.

10:07 start the target of Bellazzini:H
  • 10:23 problem with the guide in the blue channel during the dofpia. we try again.

10:23 run science Bellazzini H (2015B_1_H.ob)
  • seeing 0.8''-0.9''

10:38 DOFPIA e copointing for Castellano
  • Castellano didn't provide copointing ob, so we have to chose a star near the field from the available catalogue (WT10-318)
  • 10:58 once again problem with lbcrangebal; IRAf is not working; we have to quit and start it again to do the copointing procedure
  • 11:01 run Castelllano Science (Castellano_MACSJ0717_UR_1.ob)
  • seeing 0.9''-1''
  • in the first image in the red channel there is a bit of elongation, but the next images are good
  • 11:35 we stop the integration to run DOFPIA again because in the blue channel we have some elongation
  • 11:42 Castellano_MACSJ0717_UR_1.ob again
  • seeing 0.7''
  • in the blue channel still a bit of elongation, even if we run DOFPIA 20 min ago

12:06 running DOFPIA

12:11 Castellano_MACSJ0717_UR_2.ob (only 1 image)
  • seeing 0.8''

12:24 slewing to the photometric standard SA98
  • DOFPIA
  • 12:32 SA98_SBVg_Vri_Nov15.ob

12:45 V-I sky flats

13:30 Antoniucci LUCI calibrations (flats and arcs)
  • flats zJspec luci2.20151108.0060-0069
  • flats HKspec luci2.20151108.0070-0076 + luci2.20151108.0079-0081
  • arcs HKspec: neon luci2.20151108.0082-0083, argon luci2.20151108.0084-0085, xenon luci2.20151108.0086-0087
  • arcs zJspec: neon luci2.20151108.0088-0089 (dit=60s) luci2.20151108.0090-0091 (dit=30s), argon luci2.20151108.0092-0093

* WARNING!!! Luci2 data are not saved in /newdata but in /home/redout2/DATA/20151108 (/home/redout2/DATA/UTDATE)

-- %USERSIG{RobertaCarini - 2015-11-08}%

Comments

 
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback